# AcademicBrowse: Benchmarking Academic Browse Ability of LLMs

### v1

@@ Abstract
Large Language Models (LLMs)’ search capabilities have garnered significant attention. Existing benchmarks, such as OpenAI’s BrowseComp, primarily focus on general search scenarios and fail to adequately address the specific demands of academic search. These demands include deeper literature tracing and organization, professional support for academic databases, the ability to navigate long-tail academic knowledge, and ensuring academic rigor. Here, we proposed AcademicBrowse, the first dataset specifically designed to evaluate the complex information retrieval capabilities of Large Language Models (LLMs) in academic research. AcademicBrowse possesses the following key characteristics: Academic Practicality, where question content closely mirrors real academic learning and research environments, avoiding deliberately misleading models; High Difficulty, with answers that are challenging for single models (e.g., Grok DeepSearch or Gemini Deep Research) to provide directly, often requiring at least three deep searches to derive; Concise Evaluation, where limiting conditions ensure answers are as unique as possible, accompanied by clear sources and brief solution explanations, greatly facilitating subsequent audit and verification, surpassing the current lack of analyzed search datasets both domestically and internationally; and Broad Coverage, as the dataset spans at least 15 different academic disciplines. Through AcademicBrowse, we expect to more precisely measure and promote the performance improvement of LLMs in complex academic information retrieval tasks. The data is available at: https://huggingface.co/datasets/PKU-DS-LAB/AcademicBrowse
Figure 1: The data collection pipeline of AcademicBrowse. Data sourced by students from academic publications or websites is first filtered using several LLMs to select for difficult questions. A specialized team then reviews this selected data for uniqueness, source accessibility, and academic correctness.
@@ Introduction
To address this significant gap, we propose AcademicBrowse, a novel dataset consisting of 223 high quality questions meticulously designed to rigorously evaluate LLMs’ complex information retrieval skills in an academic context. Our dataset is distinctively collected through a unique methodology: questions are crafted such that their answers cannot be directly derived from a single search query or existing LLM knowledge, demanding iterative, multi-hop searches for resolution. Furthermore, we mandate the use of powerful search-augmented LLMs, specifically Grok DeepSearch and Gemini Deep Research, during the answer acquisition phase, ensuring that the questions posed are genuinely challenging and cannot be easily solved by current state-of-the-art models. This rigorous collection process, combined with precise answer verification and source attribution, ensures that AcademicBrowse effectively captures and quantifies an LLM’s capacity for deep, persistent, and verifiable academic information retrieval.
We are delighted to propose AcademicBrowse, a pioneering dataset that offers a comprehensive and challenging framework for evaluating LLM capabilities in complex academic information retrieval. Through its unique design and stringent data collection methodology, AcademicBrowse provides crucial insights into the strengths and weaknesses of current large language models when confronted with the intricate demands of scholarly research. We believe this dataset will serve as an invaluable resource for driving future advancements in building more capable and reliable AI systems for academic applications.
@@
Statistics of AcademicBrowse
The AcademicBrowse dataset comprises 223 meticulously curated questions, with disciplinary distribution shown in Figure 2. Each entry contains four structured components: question, answer, explanation, and domain (see Figure 3).
@@
Figure 2: Distribution of Disciplines in AcademicBrowse. The discipline for each question was manually annotated by the question creators.
Figure 3: Example of AcademicBrowse. Each data in the dataset includes a question, answer, explanation, and corresponding domain.
Testing Inference Accuracy on AcademicBrowse
The experiment evaluated several Large Language Models (LLMs) on the AcademicBrowse dataset, focusing on their inherent reasoning abilities versus the performance enhancements provided by real-time browsing. Our evaluation follows the same methodology as Humanity’s Last Exam (Phan and et.al 2025), where an LLM is used as a judge to assess the model’s output after inference. The exact prompt used for judgment is provided in AppendixEvaluation Prompt. The key findings, as outlined in Table 1, highlight the critical role that browsing capabilities play in improving the accuracy of LLMs in complex academic information retrieval tasks.
@@ Discussion
Question from AcademicBrowse Cannot Be Solved by Reasoning Alone
The first major takeaway from our experiment is that the AcademicBrowse dataset cannot be solved by reasoning alone, without the aid of external resources.
@@
This indicates that the questions within AcademicBrowse are highly complex, requiring more than just deep reasoning or inferences drawn from static knowledge. These types of questions often involve the need for up-to-date and specific information, as well as multi-hop reasoning across multiple sources. This highlights the limitation of relying solely on the knowledge embedded within pre-trained models. Academic research queries often require dynamic, cross-referenced, and sometimes real-time data that pure reasoning models cannot effectively provide.
@@
These findings underline the critical importance of research-focused models for solving real-world academic problems. AcademicBrowse serves as a valuable benchmark because it reveals the current gaps in search-based LLMs and highlights the need for more advanced research models. It confirms that a true research model must go beyond simply retrieving information, but also incorporate sophisticated techniques for synthesis, understanding context, and ensuring the accuracy of the answers. This makes AcademicBrowse a crucial tool for benchmarking models that are meant to handle academic tasks, as it provides a realistic challenge for evaluating their ability to engage in real research workflows.
@@ Conclusion
In this work, we introduced AcademicBrowse, the first meticulously curated benchmark dataset designed to specifically evaluate the complex information retrieval capabilities of Large Language Models (LLMs) in academic research settings. Unlike existing benchmarks, AcademicBrowse emphasizes the need for multi-hop, deep searches, demanding real-time information acquisition beyond static knowledge bases. Our rigorous data collection methodology, involving expert human curation and stringent LLM-based filtering, ensures the dataset’s academic practicality, high difficulty, concise evaluability, and broad disciplinary coverage.
Our experimental findings reveal several critical insights. First, we demonstrate unequivocally that academic research queries, as represented in AcademicBrowse, cannot be solved by reasoning alone, highlighting the inherent limitations of models without external Browse capabilities. Second, we empirically validated the significant performance improvement when LLMs are equipped with Browse capabilities, underscoring the necessity of real-time information retrieval for academic tasks. Despite this improvement, our results also underscore a crucial challenge: even state-of-the-art search-enabled models remain insufficient for consistently addressing complex academic problems, achieving modest accuracy. This suggests that current LLMs, while capable of basic Browse, lack advanced synthesis, contextual understanding, and accuracy verification mechanisms crucial for true academic research.
AcademicBrowse thus serves as a vital tool, providing a realistic and challenging benchmark that reveals the current gaps in search-augmented LLMs. We believe that this dataset will be instrumental in driving future research towards developing more sophisticated and reliable AI systems capable of robust, verifiable, and in-depth academic information retrieval, ultimately empowering researchers with more effective AI assistants. Future work will focus on expanding the scale and diversity of AcademicBrowse, exploring multimodal academic search, and developing more nuanced evaluation metrics that capture the quality of synthesized information and source attribution.
@@ Acknowledgments
Examples of AcademicBrowse
Figure 5: Example 1 of AcademicBrowse dataset
Figure 6: Example 2 of AcademicBrowse dataset
Figure 7: Example 3 of AcademicBrowse dataset

### v2

# ScholarSearch: Benchmarking Scholar Searching Ability of LLMs

@@ Abstract
Large Language Models (LLMs)’ search capabilities have garnered significant attention. Existing benchmarks, such as OpenAI’s BrowseComp, primarily focus on general search scenarios and fail to adequately address the specific demands of academic search. These demands include deeper literature tracing and organization, professional support for academic databases, the ability to navigate long-tail academic knowledge, and ensuring academic rigor. Here, we proposed ScholarSearch, the first dataset specifically designed to evaluate the complex information retrieval capabilities of Large Language Models (LLMs) in academic research. ScholarSearch possesses the following key characteristics: Academic Practicality, where question content closely mirrors real academic learning and research environments, avoiding deliberately misleading models; High Difficulty, with answers that are challenging for single models (e.g., Grok DeepSearch or Gemini Deep Research) to provide directly, often requiring at least three deep searches to derive; Concise Evaluation, where limiting conditions ensure answers are as unique as possible, accompanied by clear sources and brief solution explanations, greatly facilitating subsequent audit and verification, surpassing the current lack of analyzed search datasets both domestically and internationally; and Broad Coverage, as the dataset spans at least 15 different academic disciplines. Through ScholarSearch, we expect to more precisely measure and promote the performance improvement of LLMs in complex academic information retrieval tasks. The data is available at: https://huggingface.co/datasets/PKU-DS-LAB/ScholarSearch
Figure 1: The data collection pipeline of ScholarSearch. Data sourced by students from academic publications or websites is first filtered using several LLMs to select for difficult questions. A specialized team then reviews this selected data for uniqueness, source accessibility, and academic correctness.
@@ Introduction
To address this significant gap, we propose ScholarSearch, a novel dataset consisting of 223 high-quality questions meticulously designed to rigorously evaluate LLMs’ complex information retrieval skills in an academic context. Our dataset is distinctively collected through a unique methodology: questions are crafted such that their answers cannot be directly derived from a single search query or existing LLM knowledge, demanding iterative, multi-hop searches for resolution. Furthermore, we mandate the use of powerful search-augmented LLMs, specifically Grok DeepSearch and Gemini Deep Research, during the answer acquisition phase, ensuring that the questions posed are genuinely challenging and cannot be easily solved by current state-of-the-art models. This rigorous collection process, combined with precise answer verification and source attribution, ensures that ScholarSearch effectively captures and quantifies an LLM’s capacity for deep, persistent, and verifiable academic information retrieval.
We are delighted to propose ScholarSearch, a pioneering dataset that offers a comprehensive and challenging framework for evaluating LLM capabilities in complex academic information retrieval. Through its unique design and stringent data collection methodology, ScholarSearch provides crucial insights into the strengths and weaknesses of current large language models when confronted with the intricate demands of scholarly research. We believe this dataset will serve as an invaluable resource for driving future advancements in building more capable and reliable AI systems for academic applications.
@@
Statistics of ScholarSearch
The ScholarSearch dataset comprises 223 meticulously curated questions, with disciplinary distribution shown in Figure 2. Each entry contains four structured components: question, answer, explanation, and domain (see Figure 3).
@@
Figure 2: Distribution of Disciplines in ScholarSearch. The discipline for each question was manually annotated by the question creators.
Figure 3: Example of ScholarSearch. Each data in the dataset includes a question, answer, explanation, and corresponding domain.
Testing Inference Accuracy on ScholarSearch
The experiment evaluated several Large Language Models (LLMs) on the ScholarSearch dataset, focusing on their inherent reasoning abilities versus the performance enhancements provided by real-time browsing. Our evaluation follows the same methodology as Humanity’s Last Exam (Phan and et.al 2025), where an LLM is used as a judge to assess the model’s output after inference. The exact prompt used for judgment is provided in AppendixEvaluation Prompt. The key findings, as outlined in Table 1, highlight the critical role that browsing capabilities play in improving the accuracy of LLMs in complex academic information retrieval tasks.
@@ Discussion
Question from ScholarSearch Cannot Be Solved by Reasoning Alone
The first major takeaway from our experiment is that the ScholarSearch dataset cannot be solved by reasoning alone, without the aid of external resources.
@@
This indicates that the questions within ScholarSearch are highly complex, requiring more than just deep reasoning or inferences drawn from static knowledge. These types of questions often involve the need for up-to-date and specific information, as well as multi-hop reasoning across multiple sources. This highlights the limitation of relying solely on the knowledge embedded within pre-trained models. Academic research queries often require dynamic, cross-referenced, and sometimes real-time data that pure reasoning models cannot effectively provide.
@@
These findings underline the critical importance of research-focused models for solving real-world academic problems. ScholarSearch serves as a valuable benchmark because it reveals the current gaps in search-based LLMs and highlights the need for more advanced research models. It confirms that a true research model must go beyond simply retrieving information, but also incorporate sophisticated techniques for synthesis, understanding context, and ensuring the accuracy of the answers. This makes ScholarSearch a crucial tool for benchmarking models that are meant to handle academic tasks, as it provides a realistic challenge for evaluating their ability to engage in real research workflows.
@@ Conclusion
In this work, we introduced ScholarSearch, the first meticulously curated benchmark dataset designed to specifically evaluate the complex information retrieval capabilities of Large Language Models (LLMs) in academic research settings. Unlike existing benchmarks, ScholarSearch emphasizes the need for multi-hop, deep searches, demanding real-time information acquisition beyond static knowledge bases. Our rigorous data collection methodology, involving expert human curation and stringent LLM-based filtering, ensures the dataset’s academic practicality, high difficulty, concise evaluability, and broad disciplinary coverage.
Our experimental findings reveal several critical insights. First, we demonstrate unequivocally that academic research queries, as represented in ScholarSearch, cannot be solved by reasoning alone, highlighting the inherent limitations of models without external Browse capabilities. Second, we empirically validated the significant performance improvement when LLMs are equipped with Browse capabilities, underscoring the necessity of real-time information retrieval for academic tasks. Despite this improvement, our results also underscore a crucial challenge: even state-of-the-art search-enabled models remain insufficient for consistently addressing complex academic problems, achieving modest accuracy. This suggests that current LLMs, while capable of basic Browse, lack advanced synthesis, contextual understanding, and accuracy verification mechanisms crucial for true academic research.
ScholarSearch thus serves as a vital tool, providing a realistic and challenging benchmark that reveals the current gaps in search-augmented LLMs. We believe that this dataset will be instrumental in driving future research towards developing more sophisticated and reliable AI systems capable of robust, verifiable, and in-depth academic information retrieval, ultimately empowering researchers with more effective AI assistants. Future work will focus on expanding the scale and diversity of ScholarSearch, exploring multimodal academic search, and developing more nuanced evaluation metrics that capture the quality of synthesized information and source attribution.
@@ Acknowledgments
Examples of ScholarSearch
Figure 5: Example 1 of ScholarSearch dataset
Figure 6: Example 2 of ScholarSearch dataset
Figure 7: Example 3 of ScholarSearch dataset